\begin{thebibliography}{29}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bravyi et~al.(2024)Bravyi, Cross, Gambetta, Maslov, Rall, and Yoder]{Bravyi2024}
Sergey Bravyi, Andrew~W. Cross, Jay~M. Gambetta, Dmitri Maslov, Patrick Rall, and Theodore~J. Yoder.
\newblock High-threshold and low-overhead fault-tolerant quantum memory.
\newblock \emph{Nature}, 627\penalty0 (8005):\penalty0 778–782, March 2024.
\newblock ISSN 1476-4687.
\newblock \doi{10.1038/s41586-024-07107-7}.
\newblock URL \url{http://dx.doi.org/10.1038/s41586-024-07107-7}.

\bibitem[Bravyi and Kitaev(1998)]{bravyi1998quantumcodeslatticeboundary}
S.~B. Bravyi and A.~Yu. Kitaev.
\newblock Quantum codes on a lattice with boundary, 1998.
\newblock URL \url{https://arxiv.org/abs/quant-ph/9811052}.

\bibitem[Dennis et~al.(2002)Dennis, Kitaev, Landahl, and Preskill]{Dennis_2002}
Eric Dennis, Alexei Kitaev, Andrew Landahl, and John Preskill.
\newblock Topological quantum memory.
\newblock \emph{Journal of Mathematical Physics}, 43\penalty0 (9):\penalty0 4452–4505, September 2002.
\newblock ISSN 1089-7658.
\newblock \doi{10.1063/1.1499754}.
\newblock URL \url{http://dx.doi.org/10.1063/1.1499754}.

\bibitem[Acharya et~al.(2023)Acharya, Aleiner, Allen, Andersen, Ansmann, Arute, Arya, Asfaw, Atalaya, Babbush, Bacon, Bardin, Basso, Bengtsson, Boixo, Bortoli, Bourassa, Bovaird, Brill, Broughton, Buckley, Buell, Burger, Burkett, Bushnell, Chen, Chen, Chiaro, Cogan, Collins, Conner, Courtney, Crook, Curtin, Debroy, Del Toro~Barba, Demura, Dunsworth, Eppens, Erickson, Faoro, Farhi, Fatemi, Flores~Burgos, Forati, Fowler, Foxen, Giang, Gidney, Gilboa, Giustina, Grajales~Dau, Gross, Habegger, Hamilton, Harrigan, Harrington, Higgott, Hilton, Hoffmann, Hong, Huang, Huff, Huggins, Ioffe, Isakov, Iveland, Jeffrey, Jiang, Jones, Juhas, Kafri, Kechedzhi, Kelly, Khattar, Khezri, Kieferová, Kim, Kitaev, Klimov, Klots, Korotkov, Kostritsa, Kreikebaum, Landhuis, Laptev, Lau, Laws, Lee, Lee, Lester, Lill, Liu, Locharla, Lucero, Malone, Marshall, Martin, McClean, McCourt, McEwen, Megrant, Meurer~Costa, Mi, Miao, Mohseni, Montazeri, Morvan, Mount, Mruczkiewicz, Naaman, Neeley, Neill, Nersisyan, Neven, Newman, Ng, Nguyen,
  Nguyen, Niu, O’Brien, Opremcak, Platt, Petukhov, Potter, Pryadko, Quintana, Roushan, Rubin, Saei, Sank, Sankaragomathi, Satzinger, Schurkus, Schuster, Shearn, Shorter, Shvarts, Skruzny, Smelyanskiy, Smith, Sterling, Strain, Szalay, Torres, Vidal, Villalonga, Vollgraff~Heidweiller, White, Xing, Yao, Yeh, Yoo, Young, Zalcman, Zhang, and Zhu]{Google2023}
Rajeev Acharya, Igor Aleiner, Richard Allen, Trond~I. Andersen, Markus Ansmann, Frank Arute, Kunal Arya, Abraham Asfaw, Juan Atalaya, Ryan Babbush, Dave Bacon, Joseph~C. Bardin, Joao Basso, Andreas Bengtsson, Sergio Boixo, Gina Bortoli, Alexandre Bourassa, Jenna Bovaird, Leon Brill, Michael Broughton, Bob~B. Buckley, David~A. Buell, Tim Burger, Brian Burkett, Nicholas Bushnell, Yu~Chen, Zijun Chen, Ben Chiaro, Josh Cogan, Roberto Collins, Paul Conner, William Courtney, Alexander~L. Crook, Ben Curtin, Dripto~M. Debroy, Alexander Del Toro~Barba, Sean Demura, Andrew Dunsworth, Daniel Eppens, Catherine Erickson, Lara Faoro, Edward Farhi, Reza Fatemi, Leslie Flores~Burgos, Ebrahim Forati, Austin~G. Fowler, Brooks Foxen, William Giang, Craig Gidney, Dar Gilboa, Marissa Giustina, Alejandro Grajales~Dau, Jonathan~A. Gross, Steve Habegger, Michael~C. Hamilton, Matthew~P. Harrigan, Sean~D. Harrington, Oscar Higgott, Jeremy Hilton, Markus Hoffmann, Sabrina Hong, Trent Huang, Ashley Huff, William~J. Huggins, Lev~B.
  Ioffe, Sergei~V. Isakov, Justin Iveland, Evan Jeffrey, Zhang Jiang, Cody Jones, Pavol Juhas, Dvir Kafri, Kostyantyn Kechedzhi, Julian Kelly, Tanuj Khattar, Mostafa Khezri, Mária Kieferová, Seon Kim, Alexei Kitaev, Paul~V. Klimov, Andrey~R. Klots, Alexander~N. Korotkov, Fedor Kostritsa, John~Mark Kreikebaum, David Landhuis, Pavel Laptev, Kim-Ming Lau, Lily Laws, Joonho Lee, Kenny Lee, Brian~J. Lester, Alexander Lill, Wayne Liu, Aditya Locharla, Erik Lucero, Fionn~D. Malone, Jeffrey Marshall, Orion Martin, Jarrod~R. McClean, Trevor McCourt, Matt McEwen, Anthony Megrant, Bernardo Meurer~Costa, Xiao Mi, Kevin~C. Miao, Masoud Mohseni, Shirin Montazeri, Alexis Morvan, Emily Mount, Wojciech Mruczkiewicz, Ofer Naaman, Matthew Neeley, Charles Neill, Ani Nersisyan, Hartmut Neven, Michael Newman, Jiun~How Ng, Anthony Nguyen, Murray Nguyen, Murphy~Yuezhen Niu, Thomas~E. O’Brien, Alex Opremcak, John Platt, Andre Petukhov, Rebecca Potter, Leonid~P. Pryadko, Chris Quintana, Pedram Roushan, Nicholas~C. Rubin, Negar
  Saei, Daniel Sank, Kannan Sankaragomathi, Kevin~J. Satzinger, Henry~F. Schurkus, Christopher Schuster, Michael~J. Shearn, Aaron Shorter, Vladimir Shvarts, Jindra Skruzny, Vadim Smelyanskiy, W.~Clarke Smith, George Sterling, Doug Strain, Marco Szalay, Alfredo Torres, Guifre Vidal, Benjamin Villalonga, Catherine Vollgraff~Heidweiller, Theodore White, Cheng Xing, Z.~Jamie Yao, Ping Yeh, Juhwan Yoo, Grayson Young, Adam Zalcman, Yaxing Zhang, and Ningfeng Zhu.
\newblock Suppressing quantum errors by scaling a surface code logical qubit.
\newblock \emph{Nature}, 614\penalty0 (7949):\penalty0 676–681, February 2023.
\newblock ISSN 1476-4687.
\newblock \doi{10.1038/s41586-022-05434-1}.
\newblock URL \url{http://dx.doi.org/10.1038/s41586-022-05434-1}.

\bibitem[Acharya et~al.(2024)Acharya, Abanin, Aghababaie-Beni, Aleiner, Andersen, Ansmann, Arute, Arya, Asfaw, Astrakhantsev, Atalaya, Babbush, Bacon, Ballard, Bardin, Bausch, Bengtsson, Bilmes, Blackwell, Boixo, Bortoli, Bourassa, Bovaird, Brill, Broughton, Browne, Buchea, Buckley, Buell, Burger, Burkett, Bushnell, Cabrera, Campero, Chang, Chen, Chen, Chiaro, Chik, Chou, Claes, Cleland, Cogan, Collins, Conner, Courtney, Crook, Curtin, Das, Davies, De~Lorenzo, Debroy, Demura, Devoret, Di~Paolo, Donohoe, Drozdov, Dunsworth, Earle, Edlich, Eickbusch, Elbag, Elzouka, Erickson, Faoro, Farhi, Ferreira, Burgos, Forati, Fowler, Foxen, Ganjam, Garcia, Gasca, Genois, Giang, Gidney, Gilboa, Gosula, Dau, Graumann, Greene, Gross, Habegger, Hall, Hamilton, Hansen, Harrigan, Harrington, Heras, Heslin, Heu, Higgott, Hill, Hilton, Holland, Hong, Huang, Huff, Huggins, Ioffe, Isakov, Iveland, Jeffrey, Jiang, Jones, Jordan, Joshi, Juhas, Kafri, Kang, Karamlou, Kechedzhi, Kelly, Khaire, Khattar, Khezri, Kim, Klimov, Klots,
  Kobrin, Kohli, Korotkov, Kostritsa, Kothari, Kozlovskii, Kreikebaum, Kurilovich, Lacroix, Landhuis, Lange-Dei, Langley, Laptev, Lau, Le~Guevel, Ledford, Lee, Lee, Lensky, Leon, Lester, Li, Li, Lill, Liu, Livingston, Locharla, Lucero, Lundahl, Lunt, Madhuk, Malone, Maloney, Mandrà, Manyika, Martin, Martin, Martin, Maxfield, McClean, McEwen, Meeks, Megrant, Mi, Miao, Mieszala, Molavi, Molina, Montazeri, Morvan, Movassagh, Mruczkiewicz, Naaman, Neeley, Neill, Nersisyan, Neven, Newman, Ng, Nguyen, Nguyen, Ni, Niu, O’Brien, Oliver, Opremcak, Ottosson, Petukhov, Pizzuto, Platt, Potter, Pritchard, Pryadko, Quintana, Ramachandran, Reagor, Redding, Rhodes, Roberts, Rosenberg, Rosenfeld, Roushan, Rubin, Saei, Sank, Sankaragomathi, Satzinger, Schurkus, Schuster, Senior, Shearn, Shorter, Shutty, Shvarts, Singh, Sivak, Skruzny, Small, Smelyanskiy, Smith, Somma, Springer, Sterling, Strain, Suchard, Szasz, Sztein, Thor, Torres, Torunbalci, Vaishnav, Vargas, Vdovichev, Vidal, Villalonga, Heidweiller, Waltman, Wang,
  Ware, Weber, Weidel, White, Wong, Woo, Xing, Yao, Yeh, Ying, Yoo, Yosri, Young, Zalcman, Zhang, Zhu, and Zobrist]{GoogleSurface2024}
Rajeev Acharya, Dmitry~A. Abanin, Laleh Aghababaie-Beni, Igor Aleiner, Trond~I. Andersen, Markus Ansmann, Frank Arute, Kunal Arya, Abraham Asfaw, Nikita Astrakhantsev, Juan Atalaya, Ryan Babbush, Dave Bacon, Brian Ballard, Joseph~C. Bardin, Johannes Bausch, Andreas Bengtsson, Alexander Bilmes, Sam Blackwell, Sergio Boixo, Gina Bortoli, Alexandre Bourassa, Jenna Bovaird, Leon Brill, Michael Broughton, David~A. Browne, Brett Buchea, Bob~B. Buckley, David~A. Buell, Tim Burger, Brian Burkett, Nicholas Bushnell, Anthony Cabrera, Juan Campero, Hung-Shen Chang, Yu~Chen, Zijun Chen, Ben Chiaro, Desmond Chik, Charina Chou, Jahan Claes, Agnetta~Y. Cleland, Josh Cogan, Roberto Collins, Paul Conner, William Courtney, Alexander~L. Crook, Ben Curtin, Sayan Das, Alex Davies, Laura De~Lorenzo, Dripto~M. Debroy, Sean Demura, Michel Devoret, Agustin Di~Paolo, Paul Donohoe, Ilya Drozdov, Andrew Dunsworth, Clint Earle, Thomas Edlich, Alec Eickbusch, Aviv~Moshe Elbag, Mahmoud Elzouka, Catherine Erickson, Lara Faoro, Edward Farhi,
  Vinicius~S. Ferreira, Leslie~Flores Burgos, Ebrahim Forati, Austin~G. Fowler, Brooks Foxen, Suhas Ganjam, Gonzalo Garcia, Robert Gasca, Élie Genois, William Giang, Craig Gidney, Dar Gilboa, Raja Gosula, Alejandro~Grajales Dau, Dietrich Graumann, Alex Greene, Jonathan~A. Gross, Steve Habegger, John Hall, Michael~C. Hamilton, Monica Hansen, Matthew~P. Harrigan, Sean~D. Harrington, Francisco J.~H. Heras, Stephen Heslin, Paula Heu, Oscar Higgott, Gordon Hill, Jeremy Hilton, George Holland, Sabrina Hong, Hsin-Yuan Huang, Ashley Huff, William~J. Huggins, Lev~B. Ioffe, Sergei~V. Isakov, Justin Iveland, Evan Jeffrey, Zhang Jiang, Cody Jones, Stephen Jordan, Chaitali Joshi, Pavol Juhas, Dvir Kafri, Hui Kang, Amir~H. Karamlou, Kostyantyn Kechedzhi, Julian Kelly, Trupti Khaire, Tanuj Khattar, Mostafa Khezri, Seon Kim, Paul~V. Klimov, Andrey~R. Klots, Bryce Kobrin, Pushmeet Kohli, Alexander~N. Korotkov, Fedor Kostritsa, Robin Kothari, Borislav Kozlovskii, John~Mark Kreikebaum, Vladislav~D. Kurilovich, Nathan Lacroix,
  David Landhuis, Tiano Lange-Dei, Brandon~W. Langley, Pavel Laptev, Kim-Ming Lau, Loïck Le~Guevel, Justin Ledford, Joonho Lee, Kenny Lee, Yuri~D. Lensky, Shannon Leon, Brian~J. Lester, Wing~Yan Li, Yin Li, Alexander~T. Lill, Wayne Liu, William~P. Livingston, Aditya Locharla, Erik Lucero, Daniel Lundahl, Aaron Lunt, Sid Madhuk, Fionn~D. Malone, Ashley Maloney, Salvatore Mandrà, James Manyika, Leigh~S. Martin, Orion Martin, Steven Martin, Cameron Maxfield, Jarrod~R. McClean, Matt McEwen, Seneca Meeks, Anthony Megrant, Xiao Mi, Kevin~C. Miao, Amanda Mieszala, Reza Molavi, Sebastian Molina, Shirin Montazeri, Alexis Morvan, Ramis Movassagh, Wojciech Mruczkiewicz, Ofer Naaman, Matthew Neeley, Charles Neill, Ani Nersisyan, Hartmut Neven, Michael Newman, Jiun~How Ng, Anthony Nguyen, Murray Nguyen, Chia-Hung Ni, Murphy~Yuezhen Niu, Thomas~E. O’Brien, William~D. Oliver, Alex Opremcak, Kristoffer Ottosson, Andre Petukhov, Alex Pizzuto, John Platt, Rebecca Potter, Orion Pritchard, Leonid~P. Pryadko, Chris Quintana,
  Ganesh Ramachandran, Matthew~J. Reagor, John Redding, David~M. Rhodes, Gabrielle Roberts, Eliott Rosenberg, Emma Rosenfeld, Pedram Roushan, Nicholas~C. Rubin, Negar Saei, Daniel Sank, Kannan Sankaragomathi, Kevin~J. Satzinger, Henry~F. Schurkus, Christopher Schuster, Andrew~W. Senior, Michael~J. Shearn, Aaron Shorter, Noah Shutty, Vladimir Shvarts, Shraddha Singh, Volodymyr Sivak, Jindra Skruzny, Spencer Small, Vadim Smelyanskiy, W.~Clarke Smith, Rolando~D. Somma, Sofia Springer, George Sterling, Doug Strain, Jordan Suchard, Aaron Szasz, Alex Sztein, Douglas Thor, Alfredo Torres, M.~Mert Torunbalci, Abeer Vaishnav, Justin Vargas, Sergey Vdovichev, Guifre Vidal, Benjamin Villalonga, Catherine~Vollgraff Heidweiller, Steven Waltman, Shannon~X. Wang, Brayden Ware, Kate Weber, Travis Weidel, Theodore White, Kristi Wong, Bryan W.~K. Woo, Cheng Xing, Z.~Jamie Yao, Ping Yeh, Bicheng Ying, Juhwan Yoo, Noureldin Yosri, Grayson Young, Adam Zalcman, Yaxing Zhang, Ningfeng Zhu, and Nicholas Zobrist.
\newblock Quantum error correction below the surface code threshold.
\newblock \emph{Nature}, 638\penalty0 (8052):\penalty0 920–926, December 2024.
\newblock ISSN 1476-4687.
\newblock \doi{10.1038/s41586-024-08449-y}.
\newblock URL \url{http://dx.doi.org/10.1038/s41586-024-08449-y}.

\bibitem[Gidney and Bacon(2023)]{gidney2023baconthreshold}
Craig Gidney and Dave Bacon.
\newblock Less bacon more threshold, 2023.
\newblock URL \url{https://arxiv.org/abs/2305.12046}.

\bibitem[Calderbank and Shor(1996)]{Shor95}
A.~R. Calderbank and Peter~W. Shor.
\newblock Good quantum error-correcting codes exist.
\newblock \emph{Phys. Rev. A}, 54:\penalty0 1098--1105, Aug 1996.
\newblock \doi{10.1103/PhysRevA.54.1098}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevA.54.1098}.

\bibitem[Nguyen et~al.(2021)Nguyen, Li, Green, Huerta~Alderete, Zhu, Zhu, Brown, and Linke]{Nguyen_2021}
Nhung~H. Nguyen, Muyuan Li, Alaina~M. Green, C.~Huerta~Alderete, Yingyue Zhu, Daiwei Zhu, Kenneth~R. Brown, and Norbert~M. Linke.
\newblock Demonstration of shor encoding on a trapped-ion quantum computer.
\newblock \emph{Physical Review Applied}, 16\penalty0 (2), August 2021.
\newblock ISSN 2331-7019.
\newblock \doi{10.1103/physrevapplied.16.024057}.
\newblock URL \url{http://dx.doi.org/10.1103/PhysRevApplied.16.024057}.

\bibitem[Wang et~al.(2025)Wang, Lu, Zhang, Liu, Chen, Wang, Wu, Xu, Zhu, Jin, Gao, Tan, Cui, Wang, Zou, Zhang, Li, Shen, Zhong, Bao, Zhu, Han, He, Shen, Wang, Yang, Song, Deng, Dong, Sun, Li, Ye, Jiang, Ma, Shen, Zhang, Li, Guo, Wang, Song, Wang, and Deng]{wang2025demonstrationlowoverheadquantumerror}
Ke~Wang, Zhide Lu, Chuanyu Zhang, Gongyu Liu, Jiachen Chen, Yanzhe Wang, Yaozu Wu, Shibo Xu, Xuhao Zhu, Feitong Jin, Yu~Gao, Ziqi Tan, Zhengyi Cui, Ning Wang, Yiren Zou, Aosai Zhang, Tingting Li, Fanhao Shen, Jiarun Zhong, Zehang Bao, Zitian Zhu, Yihang Han, Yiyang He, Jiayuan Shen, Han Wang, Jia-Nan Yang, Zixuan Song, Jinfeng Deng, Hang Dong, Zheng-Zhi Sun, Weikang Li, Qi~Ye, Si~Jiang, Yixuan Ma, Pei-Xin Shen, Pengfei Zhang, Hekang Li, Qiujiang Guo, Zhen Wang, Chao Song, H.~Wang, and Dong-Ling Deng.
\newblock Demonstration of low-overhead quantum error correction codes, 2025.
\newblock URL \url{https://arxiv.org/abs/2505.09684}.

\bibitem[Kuo and Lai(2022)]{Kuo2022}
Kao-Yueh Kuo and Ching-Yi Lai.
\newblock Exploiting degeneracy in belief propagation decoding of quantum codes.
\newblock \emph{npj Quantum Information}, 8\penalty0 (1), September 2022.
\newblock ISSN 2056-6387.
\newblock \doi{10.1038/s41534-022-00623-2}.
\newblock URL \url{http://dx.doi.org/10.1038/s41534-022-00623-2}.

\bibitem[Sabo et~al.(2024)Sabo, Gunderman, Ide, Vasmer, and Dauphinais]{PRXQuantum.5.040302}
Eric Sabo, Lane~G. Gunderman, Benjamin Ide, Michael Vasmer, and Guillaume Dauphinais.
\newblock Weight-reduced stabilizer codes with lower overhead.
\newblock \emph{PRX Quantum}, 5:\penalty0 040302, Oct 2024.
\newblock \doi{10.1103/PRXQuantum.5.040302}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PRXQuantum.5.040302}.

\bibitem[Roffe et~al.(2020)Roffe, White, Burton, and Campbell]{PhysRevResearch.2.043423}
Joschka Roffe, David~R. White, Simon Burton, and Earl Campbell.
\newblock Decoding across the quantum low-density parity-check code landscape.
\newblock \emph{Phys. Rev. Res.}, 2:\penalty0 043423, Dec 2020.
\newblock \doi{10.1103/PhysRevResearch.2.043423}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevResearch.2.043423}.

\bibitem[Gong et~al.(2024)Gong, Cammerer, and Renes]{gong2024lowlatencyiterativedecodingqldpc}
Anqi Gong, Sebastian Cammerer, and Joseph~M. Renes.
\newblock Toward low-latency iterative decoding of qldpc codes under circuit-level noise, 2024.
\newblock URL \url{https://arxiv.org/abs/2403.18901}.

\bibitem[Raveendran and Vasi{\'{c}}(2021)]{Raveendran2021trappingsetsof}
Nithin Raveendran and Bane Vasi{\'{c}}.
\newblock Trapping {S}ets of {Q}uantum {LDPC} {C}odes.
\newblock \emph{{Quantum}}, 5:\penalty0 562, October 2021.
\newblock ISSN 2521-327X.
\newblock \doi{10.22331/q-2021-10-14-562}.
\newblock URL \url{https://doi.org/10.22331/q-2021-10-14-562}.

\bibitem[Shor(1995)]{PhysRevA.52.R2493}
Peter~W. Shor.
\newblock Scheme for reducing decoherence in quantum computer memory.
\newblock \emph{Phys. Rev. A}, 52:\penalty0 R2493--R2496, Oct 1995.
\newblock \doi{10.1103/PhysRevA.52.R2493}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevA.52.R2493}.

\bibitem[Chang and Lai(2025)]{PhysRevA.111.052602}
En-Jui Chang and Ching-Yi Lai.
\newblock High-rate amplitude-damping shor codes with immunity to collective coherent errors.
\newblock \emph{Phys. Rev. A}, 111:\penalty0 052602, May 2025.
\newblock \doi{10.1103/PhysRevA.111.052602}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevA.111.052602}.

\bibitem[Fletcher et~al.(2008)Fletcher, Shor, and Win]{4675715}
Andrew~S. Fletcher, Peter~W. Shor, and Moe~Z. Win.
\newblock Channel-adapted quantum error correction for the amplitude damping channel.
\newblock \emph{IEEE Transactions on Information Theory}, 54\penalty0 (12):\penalty0 5705--5718, 2008.
\newblock \doi{10.1109/TIT.2008.2006458}.

\bibitem[Conway and Sloane(1986)]{1057187}
J.~Conway and N.~Sloane.
\newblock Lexicographic codes: Error-correcting codes from game theory.
\newblock \emph{IEEE Transactions on Information Theory}, 32\penalty0 (3):\penalty0 337--348, 1986.
\newblock \doi{10.1109/TIT.1986.1057187}.

\bibitem[Chang(2025{\natexlab{a}})]{hwfz-c6vy}
En-Jui Chang.
\newblock High-rate extended binomial codes for multiqubit encoding.
\newblock \emph{Phys. Rev. A}, 112:\penalty0 032419, Sep 2025{\natexlab{a}}.
\newblock \doi{10.1103/hwfz-c6vy}.
\newblock URL \url{https://link.aps.org/doi/10.1103/hwfz-c6vy}.

\bibitem[Bravyi et~al.(2018)Bravyi, Englbrecht, K\"{o}nig, and Peard]{Bravyi2018}
Sergey Bravyi, Matthias Englbrecht, Robert K\"{o}nig, and Nolan Peard.
\newblock Correcting coherent errors with surface codes.
\newblock \emph{npj Quantum Information}, 4\penalty0 (1), October 2018.
\newblock ISSN 2056-6387.
\newblock \doi{10.1038/s41534-018-0106-y}.
\newblock URL \url{http://dx.doi.org/10.1038/s41534-018-0106-y}.

\bibitem[Debroy et~al.(2018)Debroy, Li, Newman, and Brown]{Debroy2018}
Dripto~M. Debroy, Muyuan Li, Michael Newman, and Kenneth~R. Brown.
\newblock Stabilizer slicing: Coherent error cancellations in low-density parity-check stabilizer codes.
\newblock \emph{Physical Review Letters}, 121\penalty0 (25), December 2018.
\newblock ISSN 1079-7114.
\newblock \doi{10.1103/physrevlett.121.250502}.
\newblock URL \url{http://dx.doi.org/10.1103/PhysRevLett.121.250502}.

\bibitem[Hu et~al.(2022)Hu, Liang, Rengaswamy, and Calderbank]{HLRC22}
Jingzhen Hu, Qingzhong Liang, Narayanan Rengaswamy, and Robert Calderbank.
\newblock Mitigating coherent noise by balancing weight-2 z-stabilizers.
\newblock \emph{IEEE Transactions on Information Theory}, 68\penalty0 (3):\penalty0 1795--1808, 2022.
\newblock \doi{10.1109/TIT.2021.3130155}.

\bibitem[M\'{a}rton and Asb\'{o}th(2023)]{Mrton2023}
\'{A}ron M\'{a}rton and J\'{a}nos~K. Asb\'{o}th.
\newblock Coherent errors and readout errors in the surface code.
\newblock \emph{Quantum}, 7:\penalty0 1116, September 2023.
\newblock ISSN 2521-327X.
\newblock \doi{10.22331/q-2023-09-21-1116}.
\newblock URL \url{http://dx.doi.org/10.22331/q-2023-09-21-1116}.

\bibitem[Ouyang(2014)]{PhysRevA.90.062317}
Yingkai Ouyang.
\newblock Permutation-invariant quantum codes.
\newblock \emph{Phys. Rev. A}, 90:\penalty0 062317, Dec 2014.
\newblock \doi{10.1103/PhysRevA.90.062317}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevA.90.062317}.

\bibitem[Ouyang and Fitzsimons(2016)]{PhysRevA.93.042340}
Yingkai Ouyang and Joseph Fitzsimons.
\newblock Permutation-invariant codes encoding more than one qubit.
\newblock \emph{Phys. Rev. A}, 93:\penalty0 042340, Apr 2016.
\newblock \doi{10.1103/PhysRevA.93.042340}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevA.93.042340}.

\bibitem[Chang(2025{\natexlab{b}})]{qzd2-b2mx}
En-Jui Chang.
\newblock Quantum dual extended hamming code immune to collective coherent errors.
\newblock \emph{Phys. Rev. A}, 112:\penalty0 052410, Nov 2025{\natexlab{b}}.
\newblock \doi{10.1103/qzd2-b2mx}.
\newblock URL \url{https://link.aps.org/doi/10.1103/qzd2-b2mx}.

\bibitem[Michael et~al.(2016)Michael, Silveri, Brierley, Albert, Salmilehto, Jiang, and Girvin]{PhysRevX.6.031006}
Marios~H. Michael, Matti Silveri, R.~T. Brierley, Victor~V. Albert, Juha Salmilehto, Liang Jiang, and S.~M. Girvin.
\newblock New class of quantum error-correcting codes for a bosonic mode.
\newblock \emph{Phys. Rev. X}, 6:\penalty0 031006, Jul 2016.
\newblock \doi{10.1103/PhysRevX.6.031006}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevX.6.031006}.

\bibitem[Cochrane et~al.(1999)Cochrane, Milburn, and Munro]{PhysRevA.59.2631}
P.~T. Cochrane, G.~J. Milburn, and W.~J. Munro.
\newblock Macroscopically distinct quantum-superposition states as a bosonic code for amplitude damping.
\newblock \emph{Phys. Rev. A}, 59:\penalty0 2631--2634, Apr 1999.
\newblock \doi{10.1103/PhysRevA.59.2631}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevA.59.2631}.

\bibitem[Gottesman et~al.(2001)Gottesman, Kitaev, and Preskill]{PhysRevA.64.012310}
Daniel Gottesman, Alexei Kitaev, and John Preskill.
\newblock Encoding a qubit in an oscillator.
\newblock \emph{Phys. Rev. A}, 64:\penalty0 012310, Jun 2001.
\newblock \doi{10.1103/PhysRevA.64.012310}.
\newblock URL \url{https://link.aps.org/doi/10.1103/PhysRevA.64.012310}.

\end{thebibliography}

\documentclass[
reprint,
superscriptaddress,
amsmath,amssymb,
aps,
pra,
floatfix,
]{revtex4-2}

\def\final{0}

\usepackage{braket}
\usepackage{mathtools}
\usepackage{graphicx}
\usepackage{color, xcolor}
\usepackage{dcolumn}
\newcolumntype{z}[1]{D{.}{.}{#1}}
\usepackage{physics}
\usepackage{upgreek}
\usepackage{subfigure}
\usepackage{tabularx}
\newcolumntype{Y}{>{\centering\arraybackslash}X}

\usepackage[linesnumbered,ruled ]{algorithm2e}
\SetKwInput{KwOutput}{Output}
\SetKwInOut{Input}{Input}
\SetKwInput{Output}{Output}
\UseRawInputEncoding 

\usepackage{xargs}
\usepackage[colorinlistoftodos,prependcaption]{todonotes}

\definecolor{Asparagus}{rgb}{0.53, 0.66, 0.42}
\definecolor{cornflowerblue}{rgb}{0.39, 0.58, 0.93}
\definecolor{darkolivegreen}{rgb}{0.33, 0.42, 0.18}
\definecolor{awesome}{rgb}{1.0, 0.13, 0.32}

\newtheorem{definitionenv}{Definition}
\newtheorem{lemmaenv}[definitionenv]{Lemma}
\newtheorem{theoremenv}[definitionenv]{Theorem}
\newtheorem{corollaryenv}[definitionenv]{Corollary}
\newtheorem{propositionenv}[definitionenv]{Proposition}
\newtheorem{conjectureenv}[definitionenv]{Conjecture}
\newtheorem{remarkenv}[definitionenv]{Remark}
\newenvironment{remark}{\begin{remarkenv}\rm}{\end{remarkenv}}
\newcommand{\br}{\begin{remark}}
	\newcommand{\er}{\end{remark}}

\newtheorem{exampleenv}{Example}
\newtheorem{app-lemmaenv}[section]{Lemma}

\newenvironment{definition}{\begin{definitionenv}\rm}{\end{definitionenv}}
\newenvironment{lemma}{\begin{lemmaenv}\rm}{\end{lemmaenv}}
\newenvironment{theorem}{\begin{theoremenv}\rm}{\end{theoremenv}}
\newenvironment{corollary}{\begin{corollaryenv}\rm}{\end{corollaryenv}}
\newenvironment{example}{\begin{exampleenv}\rm}{\end{exampleenv}}
\newenvironment{proposition}{\begin{propositionenv}\rm}{\end{propositionenv}}
\newenvironment{conjecture}{\begin{conjectureenv}\rm}{\end{conjectureenv}}
\newenvironment{app-lemma}{\begin{app-lemmaenv}\rm}{\end{app-lemmaenv}}

\newcommand{\bd}{\begin{definition}}
	\newcommand{\ed}{\end{definition}}
\newcommand{\bl}{\begin{lemma}}
	\newcommand{\el}{\end{lemma}}
\newcommand{\elp}{\hspace*{\fill} $\Box$
\end{lemma}}
\newcommand{\bt}{\begin{theorem}}
\newcommand{\et}{\end{theorem}}
\newcommand{\etp}{\hspace*{\fill} $\Box$
\end{theorem}}
\newcommand{\bc}{\begin{corollary}}
\newcommand{\ec}{\end{corollary}}
\newcommand{\ecp}{\hspace*{\fill} $\Box$
\end{corollary}}
\newcommand{\bcj}{\begin{conjecture}}
\newcommand{\ecj}{\end{conjecture}}

\newcommand{\be}{\begin{example}}
\newcommand{\ee}{\end{example}}
\newcommand{\eep}{\hspace*{\fill} $\Box$
\end{example}}
\newcommand{\bp}{\begin{proposition}}
\newcommand{\ep}{\end{proposition}}
\newcommand{\epp}{
\end{proposition}}

\ifnum\final=0
\newcommand{\mynote}[2]{{\color{#1} \marginpar{\tiny #2}}}
\newcommand{\mybignote}[2]{{\color{#1} $\langle \langle$ #2$\rangle \rangle$}}
\newcommandx{\rednote}[2][1=]{\todo[linecolor=red,backgroundcolor=red!25,bordercolor=red,#1]{#2}}
\newcommandx{\bluenote}[2][1=]{\todo[linecolor=blue,backgroundcolor=blue!25,bordercolor=blue,#1]{#2}}
\newcommandx{\yellownote}[2][1=]{\todo[linecolor=yellow,backgroundcolor=yellow!25,bordercolor=yellow,#1]{#2}}
\newcommandx{\greennote}[2][1=]{\todo[inline,linecolor=olive,backgroundcolor=green!25,bordercolor=olive,#1]{#2}}

\else
\newcommand{\mynote}[2]{}
\newcommand{\mybignote}[2]{}
\newcommand{\rednote}[2][1=]{}
\newcommand{\bluenote}[2][1=]{}
\newcommand{\greennote}[2][1=]{}
\newcommand{\yellownote}[2][1=]{}

\fi

\usepackage{adjustbox}
\usetikzlibrary{quantikz2}
\usetikzlibrary{backgrounds,fit,decorations.pathreplacing}  
\usetikzlibrary{automata} 
\usetikzlibrary{circuits.logic.CDH}
\usetikzlibrary{circuits.logic.US}
\usetikzlibrary{mindmap}
\usetikzlibrary{decorations}
\usetikzlibrary{decorations.pathmorphing}
\usetikzlibrary{arrows,decorations.pathreplacing}

\tikzset{meter/.append style={draw, inner sep=10, rectangle, font=\vphantom{A}, minimum width=30, line width=.4, path picture={\draw[black] ([shift={(.1,.3)}]path picture bounding box.south west) to[bend left=50] ([shift={(-.1,.3)}]path picture bounding box.south east);\draw[black,-latex] ([shift={(0,.1)}]path picture bounding box.south) -- ([shift={(.3,-.1)}]path picture bounding box.north);}}}

\begin{document}

\title{
Overlapped-repetition Shor codes achieving fourfold asymptotic rate
}

\author{En-Jui Chang}
\email{phyenjui@gmail.com}
\affiliation{Taichung 421786, Taiwan}

\date{\today}

\begin{abstract}
Introducing controlled overlap among a few repetition blocks yields a fourfold asymptotic rate improvement while preserving an average stabilizer weight of \(4\). Substituting the overlapped outer layer with an LDPC code further produces a family of constructions with asymptotic rate \(2/d\). We also describe a constant-excitation variant that suppresses collective coherent errors without additional overhead, as well as a bosonic generalization that extends the framework to oscillator encodings. The resulting family achieves a code rate intermediate between that of the rotated surface code (average stabilizer weight \(4\)) and the BB code (average stabilizer weight \(6\)), while remaining free of the performance degradation typical of iterative decoders.
\end{abstract}

\maketitle

\section{Introduction}

The implementation overhead of a quantum error-correcting code (QECC) is a central practical constraint, independent of the target code distance. For a device requiring a fixed distance and a prescribed number of logical qubits to support useful quantum applications, a higher code rate directly reduces the physical-qubit resources. This reduction is particularly valuable for experimental platforms operating under stringent hardware and budgetary limitations. Notably, the bivariate bicycle (BB) codes developed by the IBM Quantum team~\cite{Bravyi2024} achieve an error threshold of \(0.7\%\) while offering a code rate roughly an order of magnitude higher than that of the surface code~\cite{bravyi1998quantumcodeslatticeboundary, Dennis_2002}, which underlies recent demonstrations by Google Quantum AI~\cite{Google2023, GoogleSurface2024}.

Beyond code rate, the stabilizer weight is an equally important design parameter. As emphasized by Gidney and Bacon~\cite{gidney2023baconthreshold}, stabilizers of larger weight are more vulnerable to measurement faults. Even for the canonical Shor code~\cite{Shor95}, Nguyen \emph{et al.}~\cite{Nguyen_2021} experimentally showed that extracting two weight-\(6\) stabilizers associated with degenerate \(Z\) errors introduces more noise than measuring six weight-\(2\) stabilizers for \(X\) errors. This sensitivity partly explains the strong performance of the surface code~\cite{bravyi1998quantumcodeslatticeboundary, Dennis_2002}, whose stabilizers have weight at most four, as demonstrated in recent Google experiments~\cite{Google2023, GoogleSurface2024}. While BB codes substantially improve the code rate, they do so at the expense of increased average stabilizer weight, from about \(4\) in the surface code to \(6\) in BB codes, and current experimental results~\cite{wang2025demonstrationlowoverheadquantumerror} indicate that BB codes with larger distance have not yet realized clear logical-error suppression.

Moreover, increased stabilizer weight leads to higher decoding complexity for belief-propagation (BP) based decoders, whose cost scales linearly with the average check weight when processing syndrome data~\cite{Kuo2022} and leads more measurement errors~\cite{PRXQuantum.5.040302}. This affects both the BP + ordered statistics decoding (OSD)~\cite{PhysRevResearch.2.043423} employed by the IBM Quantum team~\cite{Bravyi2024} and the sliding-window decoder used at NVIDIA~\cite{gong2024lowlatencyiterativedecodingqldpc}, as each window invokes BP. The resulting complexity is reflected in NVIDIA's report that their implementation ``\dots achieves a worst-case decoding latency of \(3\text{ ms}\) per window for the \([[144,12,12]]\) code.'' For comparison, Google notes that ``the qubits have a mean operating \(T_1\) of \(68\,\mu\text{s}\)" and that they perform ``real-time decoding, achieving an average latency of \(63\,\mu\text{s}\)." Even allowing for hardware differences, the roughly \(47\times\) larger latency highlights a substantial performance gap and may constitute a significant bottleneck for reaching subthreshold logical error suppression.

Beyond the aforementioned concerns regarding decoding complexity, iterative decoders can also suffer performance loss when certain error syndromes cannot be resolved within practical time limits. In particular, trapping sets~\cite{Raveendran2021trappingsetsof} may prevent the decoder from identifying a valid recovery operator. For a generic (possibly random) \([[n,k,d]]\) code, the syndrome space contains \(2^{\,n-k}\) elements, and the absence of exploitable structure makes decoding increasingly difficult as this space grows exponentially. Consequently, proposed general-purpose decoders may struggle to reliably handle all syndromes in realistic scenarios.

A straightforward strategy for mitigating such performance loss is to revisit codes with highly regular structure. For example, a \([[d^{2},1,d]]\) Shor code~\cite{PhysRevA.52.R2493, Shor95} contains \(d(d-1)\) weight-\(2\) \(ZZ\) stabilizers and \((d-1)\) weight-\(2d\) \(X^{\otimes 2d}\) stabilizers, requiring only \((1+d)2^{\,d-1}\) recovery rules rather than the full \(2^{\,d^{2}-1}\) syndrome space. Structurally, the Shor code is built from an outer \([d,1,d]\) repetition code to protect against phase-flip errors and an inner \([d,1,d]\) repetition code to protect against bit-flip errors. The average stabilizer weight is \(\tfrac{4d}{d+1} \le 4\), ensuring modest measurement overhead. The drawback, however, is the substantial qubit overhead required to achieve larger distances, resulting in a poor code rate. This trade-off motivates the search for more efficient designs that improve rate while preserving these advantageous structural features.

Indeed, the outer code can be generalized to yield a high-rate amplitude-damping \(w\) code with parameters \([[(w+1)(w+k),\,k]]\)~\cite{PhysRevA.111.052602}, capable of correcting all amplitude-damping errors of weight up to \(w\). Such a construction is intrinsically error-biased. This observation suggests that the Shor code can be extended to encode multiple logical qubits simultaneously. The generalization of the outer code in Ref.~\cite{PhysRevA.111.052602} was itself inspired by the scheme of Fletcher \emph{et al.}~\cite{4675715}, where the outer code is a \([k+1,k,2]\) single-parity-check code (SPCC) that provides limited protection against phase-flip errors. Since the SPCC is a special case of the \([n,k,d=2]\) lexicodes introduced by Conway and Sloane~\cite{1057187}, it is natural to seek analogous constructions using higher-distance lexicodes that can encode multiple qubits.

The simplest nontrivial example is the \([5,2,3]\) lexicode, which may be viewed as two overlapping \([3,1,3]\) repetition codes sharing one bit. Generalizing this idea, one can form a family of codes by overlapping \(k\) \([d,1,d]\) repetition codes with \(\ell \le \lfloor d/2 \rfloor\) shared bits, producing a \([k(d-\ell)+\ell,\,k,\,d]\) code. In the limiting case \(k=1\), this reduces to the standard repetition code, while for larger \(k\) the rate improves from \(1/d\) to \(1/(d-\ell)\), allowing up to a twofold enhancement.

After introducing in Sec.~\ref{subsec:overlap} an alternative classical code to replace the repetition code used in the \([[d^{2},1,d]]\) Shor construction, we proposed in Sec.~\ref{sec:construction} three integration strategies. The optimal parameters achieved are
\begin{equation}
\bigl[\!\bigl[\,k\lceil d/2 \rceil^{2}
     + \lfloor d/2 \rfloor\bigl(\lceil d/2 \rceil + 1\bigr)
     + (d-1)\lceil d/2 \rceil,\, k,\, d\,\bigr]\!\bigr].
\end{equation}
All three strategies yield an asymptotic average stabilizer weight of \(4\). A detailed comparison with the surface code and the BB code is provided in Sec.~\ref{subsec:compare}. Some further improved constructions are discussed in Sec.~\ref{sec:further}. Finally, we conclude in Sec.~\ref{sec:conclude}.

\section{Preliminaries}

A quantum stabilizer code that encodes $k$ logical qubits into $n$ physical qubits with code distance $d$ and code rate $k/n$ is denoted by $[[n,k,d]]$. This notation parallels that of a classical linear code, where a code encoding $k$ data bits into $n$ total bits with Hamming distance $d$ is denoted by $[n,k,d]$. In both cases, the code distance corresponds to the minimum weight of an undetectable error.

To fully appreciate the meaning of code distance, it is important to distinguish between two operational settings:
(1) the \textit{forward error correction} (FEC) scenario, in which recovery must be performed solely by the receiver without feedback; and
(2) the \textit{automatic repeat request} (ARQ) scenario, where the receiver can request retransmission upon detecting an error.
In the ARQ setting, any detectable error can be effectively corrected through retransmission, whereas in the FEC setting, one typically adopts the conventional statement that ``an $[[n,k,d]]$ code can correct up to $\lfloor (d-1)/2 \rfloor$ errors and detect up to $d-1$ errors.''

To our knowledge, many students and even some instructors interpret this conventional statement in an overly classical sense: that an error of weight $(d-1)/2 < w \le d-1$ can be detected but not corrected because it is closer to another codeword. However, a more precise formulation is that an $[[n,k,d]]$ code can correct \textit{all} stochastic Pauli errors of weight $w \in {1, \dots, \lfloor (d-1)/2 \rfloor}$ and detect \textit{all} such errors of weight $w \in {1, \dots, d-1}$. Certain higher-weight errors, though not all, may also remain correctable or at least detectable, depending on the specific structure of the code.

To elucidate the structure of both the original Shor code and the proposed enhanced constructions, we first revisit the detailed formulation of the Shor code to prevent potential confusion. The identity operator $I$ and the Pauli operators $X,Y,Z$ are defined as $I = \begin{bmatrix} 1 & 0 \\ 0 & 1  \end{bmatrix}$, $X = \begin{bmatrix} 0 & 1 \\ 1 & 0  \end{bmatrix}$, $Y = \begin{bmatrix} 0 & -\mathrm{i} \\ \mathrm{i} & 0  \end{bmatrix}$, $Z = \begin{bmatrix} 1 & 0 \\ 0 & -1  \end{bmatrix},$ with $\mathrm{i}=\sqrt{-1}$. A subscript denotes the qubit on which the Pauli operator acts; for example, \( X_j \) applies the Pauli \( X \) operator to the \( j \)-th qubit, with \( j \) a non-negative integer.

\subsection{Shor code}

The $[[9,1,3]]$ Shor code is a concatenated code composed of inner and outer repetition codes. It employs two weight-6 stabilizers, \(\{g_0,g_1\}\), for degenerate $Z$ errors and six weight-2 stabilizers, \(\{g_2,\dots,g_7\}\),  for $X$ errors.
Herein, the average stabilizer weight is \(3\).

These stabilizers can be classified into one outer group and three inner groups:
\[
\{(g_0, g_1)\}, \quad \{(g_2, g_3)\}, \quad \{(g_4, g_5)\}, \quad \{(g_6, g_7)\},
\]
where each group consists of two stabilizers forming a repetition code. The first group arises from the outer code, which corrects single phase-flip errors, while the remaining three groups correspond to the inner repetition codes. Single-qubit $Z$ errors can be organized into three degenerate sets, $(Z_{i+0}, Z_{i+1}, Z_{i+2})$ for $i \in \{0, 3, 6\}$, so only two weight-6 stabilizers are required to identify such errors.

The complete syndrome table for all single-qubit errors is presented in Table~\ref{tab:syndrome_913_single}, and selected syndromes for weight-two errors are listed in Table~\ref{tab:syndrome_913_twice}. We observe that once the code can correct single $X$ and $Z$ errors, it can also correct certain weight-two $XZ$ errors, provided the error locations do not overlap. If they do overlap, the combined effect is equivalent to a single $Y$ error.

\begin{table}[ht]
\caption{Syndrome table for single-qubit errors in the $[[9,1,3]]$ Shor code. Degeneracy among $Z$ errors is reflected by identical syndromes.}
\label{tab:syndrome_913_single}
\begin{tabularx}{\linewidth}{ c  c  c  c  c  c  c  c  c }
    \hline\hline
    Error & Syndrome & \hspace{0.4cm} & Error & Syndrome & \hspace{0.4cm} & Error & Syndrome\\
    \hline
    $X_{0}$ & 00100000 & & $Y_{0}$ & 11100000 & & $Z_{0}$ & 11000000\\
    $X_{1}$ & 00110000 & & $Y_{1}$ & 11110000 & & $Z_{1}$ & 11000000\\
    $X_{2}$ & 00010000 & & $Y_{2}$ & 11010000 & & $Z_{2}$ & 11000000\\
    $X_{3}$ & 00001000 & & $Y_{3}$ & 10001000 & & $Z_{3}$ & 10000000\\
    $X_{4}$ & 00001100 & & $Y_{4}$ & 10001100 & & $Z_{4}$ & 10000000\\
    $X_{5}$ & 00000100 & & $Y_{5}$ & 10000100 & & $Z_{5}$ & 10000000\\
    $X_{6}$ & 00000010 & & $Y_{6}$ & 01000010 & & $Z_{6}$ & 01000000\\
    $X_{7}$ & 00000011 & & $Y_{7}$ & 01000011 & & $Z_{7}$ & 01000000\\
    $X_{8}$ & 00000001 & & $Y_{8}$ & 01000001 & & $Z_{8}$ & 01000000\\
    \hline\hline
\end{tabularx}
\end{table}

\begin{table}[ht]
\caption{Syndrome extraction table of weight-two error for the $[[9,1,3]]$ code. Here, we just list a few as it is sufficient to show some errors of weight higher than one is correctable.}
\label{tab:syndrome_913_twice}
\begin{tabularx}{\linewidth}{ c  c  c  c  c }
    \hline
    \hline
    Error & Syndrome & \hspace{3.5cm} & Error & Syndrome\\
    \hline
    $X_{0}Z_{3}$  & 10100000 & $\cdots$ & $Y_{0}Z_{3}$  & 01100000\\
    $X_{1}Z_{3}$  & 10110000 & $\cdots$ & $Y_{1}Z_{3}$  & 01110000\\
    $X_{2}Z_{3}$  & 10010000 & $\cdots$ & $Y_{2}Z_{3}$  & 01010000\\
    $X_{6}Z_{3}$  & 10000010 & $\cdots$ & $Y_{6}Z_{3}$  & 11000010\\
    $X_{7}Z_{3}$  & 10000011 & $\cdots$ & $Y_{7}Z_{3}$  & 11000011\\
    $X_{8}Z_{3}$  & 10000001 & $\cdots$ & $Y_{8}Z_{3}$  & 11000001\\
    \hline
    \hline
\end{tabularx}
\end{table}

The reason we emphasize that certain higher-weight errors can still be corrected or detected is that this distinction significantly influences the estimation of the error threshold at which a particular QECC becomes advantageous. While most QECC analyses treat high-weight errors as ``negligible'' perturbative contributions, largely to avoid the substantial complexity of constructing exhaustive lookup tables for recovery operations, this simplification can lead to an underestimation of a code's actual capability. In many cases, recovery can be governed by simple, rule-based operations; for example, when a specific syndrome is flagged, a corresponding correction can be directly applied. Although a full lookup table scales exponentially as $2^r$ with the number of stabilizers $r$, implementing $r$ simple, syndrome-dependent recovery rules remains tractable.

In the Shor code, the eight stabilizers can be organized into four independent two-bit groups, reducing the decoding task to \((1+3)2^{2}=16\) elementary recovery rules for the full \(2^{8}=256\) syndrome patterns rather than correcting only \(9\times 3\) weight-\(1\) errors. By contrast, treating the code as an unstructured \([[9,1,3]]\) stabilizer code and applying an iterative BP decoder incurs a decoding cost of order \(O(N W \log\!\log N)\)~\cite{Kuo2022}, with \(N=9\) physical qubits and average stabilizer weight \(W=3\).

More complex codes sacrifice this simplicity to achieve higher code rates, which can increase the decoding time, the performance loss and the risk of additional errors during recovery. Since a primary motivation for employing protected qubits is to achieve higher computational performance compared with classical algorithms, we certainly wish to avoid situations where the exponentially growing syndrome lookup table becomes a new bottleneck.

\subsection{Overlapped repetition code}\label{subsec:overlap}

In this subsection, we construct the
\begin{equation}
[k(d-\ell) + \ell,\, k,\, d]
\end{equation}
\textit{overlapped repetition code} for $\ell \le \lfloor d/2 \rfloor$. As indicated by its definition, we start with $k$ independent $[d,1,d]$ repetition codes and merge $k\ell$ qubits into $\ell$ shared qubits, thereby reducing the total number of physical qubits while preserving the overall distance.

Each logical bit-flip operation acting on a single logical qubit has weight $d$. Any operation simultaneously flipping two logical qubits has weight $2(d - \ell) \ge d$, and operations acting on more logical qubits have weight at least $d$. Consequently, the resulting construction indeed realizes a $[k(d-\ell) + \ell,\, k,\, d]$ code.

If this classical code is interpreted as a biased quantum code, the total of $(k(d - \ell - 1) + \ell)$ stabilizers can be grouped into $(k + 2)$ distinct sets. The first $k$ groups correspond to the unshared portions of the initial repetition codes, each containing $(d - \ell - 1)$ weight-\(2\) stabilizers. The next group represents the shared portion, consisting of $(\ell - 1)$ weight-\(2\) stabilizers that retain the repetition-code structure. The final group contains a single weight-$(k + 1)$ stabilizer that connects the $k$ unshared parts with the shared region.

This grouped structure implies that only $k 2^{d - \ell - 1} + 2^{\ell - 1} + 2$ pieces of information are required for decoding, rather than the full syndrome table of size $2^{k(d - \ell - 1) + \ell}$. For moderate $d$ and large $k$, the resulting decoding complexity scales approximately as the $k$-th root of the naive exhaustive approach.

\section{Quantum overlapped-repetition}\label{sec:construction}

Having established the desired classical code, we now consider its integration within the quantum code architecture, either as the outer code, as the inner code, or simultaneously in both roles. Since the construction employing it as the outer code is the most straightforward, we first present that case, followed by the inner-code and dual-application constructions.

\subsection{The outer code case}\label{subsec:outer}

When the $[k(d-\ell) + \ell,\, k,\, d]$ \textit{overlapped repetition code} is employed as the outer code, it can be straightforwardly concatenated with a $[d,1,d]$ inner repetition code, yielding a quantum code with parameters
\begin{equation}
[[d(k(d-\ell) + \ell),\, k,\, d]].
\end{equation}
In the limiting case $k = 1$, this construction reduces to the standard $[[d^2, 1, d]]$ Shor code. Conversely, as $k \rightarrow \infty$, the asymptotic code rate approaches $\tfrac{1}{d(d-\ell)}$. For even $d$ and maximal overlap $\ell = d/2$, the rate becomes $2/d^2$, exactly twice that of the original Shor code. For odd $d$, the corresponding rate is $\tfrac{2}{d(d+1)}$.

The outer layer contributes \(k(d-\ell-1)+\ell\) stabilizers composed entirely of \(X\) operators, one of weight \((k+1)d\) and the remainder of weight \(2d\). The inner layer contributes \((k(d-\ell)+\ell)(d-1)\) stabilizers composed entirely of \(Z\) operators, each of weight \(2\).
The average stabilizer weight is
\begin{equation}
    W_{\mathrm{outer}}
    = \frac{(2(d-\ell)(2d-1)-d)k + 2\ell(2d-1) - d}{((d-\ell)d - 1)k + \ell d},
\end{equation}
which approaches \(W_{\mathrm{outer}} \to 4\) in the limit \(k \gg d \gg 1\).

\subsubsection{Protection}

To verify that the concatenated construction indeed attains distance \(d\), we analyze all error patterns of weight at most \(d\).

Suppose first that \(d' \le d\) \(X\)-type errors occur within a single inner repetition block of distance \(d\). Such an event is uniquely identifiable by the inner code, and thus any distribution of up to \(d\) \(X\) errors across the blocks is detectable.

Conversely, consider \(Z\)-type errors. If \(d' \le d\) \(Z\) errors occur within a single inner repetition block, the inner-code stabilizers ensure that the error has no effect when \(d'\) is even, and induces a single phase flip when \(d'\) is odd. In the worst case, therefore, up to \(d\) such phase flips may be produced across the blocks, each corresponding to an effective \(Z\)-error on one outer-code qubit. Since the outer repetition code has distance \(d\), all such patterns remain detectable.

Taken together, these observations confirm that the concatenated construction has distance \(d\).

\subsubsection{Relation to other codes}

Although Sec.~\ref{subsec:overlap} focuses on the regime \(\ell \le \lfloor d/2 \rfloor\), the underlying idea of overlap has antecedents in other high-rate constructions. One example is the amplitude-damping \(w\) code with parameters \([[(w+1)(w+k),\,k]]\)~\cite{PhysRevA.111.052602,4675715}, which corresponds to choosing \(d = w+1\) and \(\ell = w\). A related perspective appears in bosonic encodings such as the extended binomial code~\cite{hwfz-c6vy}, where excitations within the inner repetition structure can be merged and treated as indistinguishable particles. The essential distinction is that stronger overlap becomes permissible when the physical noise is biased, enabling higher-rate designs that would otherwise violate distance constraints in the unbiased setting.

\subsection{The inner code case}\label{subsec:inner}


In this approach, the inner layer is chosen to be the
\begin{equation}
[\,((k+1)d - 1)(d-\ell) + \ell,\; (k+1)d - 1,\; d\,]
\end{equation}
overlapped repetition code. The outer layer consists of \(k+1\) independent \([d,1,d]\) repetition codes, yielding a concatenated construction with parameters
\begin{equation}
[[\,((k+1)d - 1)(d-\ell) + \ell,\; k,\; d\,]].
\end{equation}
These parameters parallel those of the outer-code variant but may increase the total qubit count by \((d-1)(d-2\ell)\) for fixed logical dimension while preserving the distance. The asymptotic code rate remains twice that of the original Shor construction.

In this setting, the inner layer contributes \((kd + (d-1))(d-\ell-1) + \ell\) stabilizers of \(Z\)-type: one of weight \((k+1)d\) and the remainder of weight \(2\). The outer layer contributes \((k+1)(d-1)\) \(X\)-type stabilizers, of which \(k(d-1)+(d-2)\) have weight \(2(d-\ell)\), and one has weight \(d\).
The average stabilizer weight is
\begin{equation}
    W_{\mathrm{inner}}
    =
    \frac{(2(d-\ell)(2d-1)-d)\,k + 2(d-\ell)(2d-5) + 2d}
         {(d^2 - 1 - d\ell)\,k + (d-1)(d-\ell)\ell},
\end{equation}
which approaches \(W_{\mathrm{inner}}\to 4\) in the limit \(k \gg d \gg 1\).

\subsubsection*{Protection}

Again, to verify that the concatenated construction indeed attains distance \(d\), we analyze all error patterns of weight at most \(d\).

Suppose first that \(d' \le d\) \(X\)-type errors occur within the inner code of distance \(d\). Such an event is uniquely identifiable by the inner code, and hence all configurations with at most \(d\) \(X\) errors are detectable.

For \(Z\)-type errors, the analysis changes because the inner layer is now a single overlapped-repetition code rather than multiple independent repetition codes. This code consists of one shared repetition block of length \(\ell\) and several unshared blocks of length \(d - \ell\). If \(d'\) \(Z\) errors occur within any single block, the effect is trivial when \(d'\) is even and corresponds to a single phase flip when \(d'\) is odd. Thus, in the worst case, up to \(d\) such phase flips may arise across the blocks. If the inner code has parameters \([[k_i(d-\ell)+\ell,\, k_i,\, d]]\), then it contains \(k_i + 1\) blocks, and the outer layer must be designed to account for this additional block.

Consider the choice \(k_i = (k+1)d - 1\). In this setting, \(k\) identical outer repetition codes of distance \(d\) each protect \(d\) blocks, and an additional outer repetition code of distance \(d\) protects the remaining \(d-1\) unshared blocks of length \(d - \ell\) together with the single shared block of length \(\ell\). Since every outer code has distance \(d\), all such error patterns remain detectable.

Taken together, these observations confirm that the concatenated construction indeed has distance \(d\).

\subsection{Both outer and inner codes}\label{subsec:both}

Finally, the overlapping strategy may be applied simultaneously to both the inner and outer layers to further enhance the code rate. As shown in Sec.~\ref{subsec:inner}, the construction requires one additional repetition code beyond the \(k\) identical outer repetition codes. Moreover, Sec.~\ref{subsec:outer} established that these \(k\) outer repetition codes can themselves be merged into a single overlapped-repetition code. Consequently, the outer layer consists of a \([\,k(d-\ell)+\ell,\, k,\, d\,]\) code together with an additional \([\,d,1,d\,]\) repetition code.

These two outer codes require the inner layer to supply \(k_i = k(d-\ell) + \ell + (d-1)\) unshared blocks, along with one shared block. Thus, the inner layer is a \([\,k_i(d-\ell)+\ell,\, k_i,\, d\,]\) overlapped-repetition code. The resulting concatenated quantum code therefore has parameters
\begin{equation}
[[\,k(d-\ell)^2 + \ell(d-\ell) + (d-1)(d-\ell) + \ell,\, k,\, d\,]].
\end{equation}
As $k \rightarrow \infty$, the asymptotic code rate approaches $1/(d-\ell)^2$. Choosing the maximal overlap $\ell = \lfloor d/2\rfloor$ yields an asymptotic rate that is about four times higher than that of the original Shor code.

Among the inner-layer stabilizers, a single stabilizer has weight \(k_i + 1\), while all remaining stabilizers have weight \(2\). For the overlapped-repetition portion of the outer layer, one stabilizer has weight
\begin{equation}
w_k =
\begin{cases}
(k+1)(d-\ell), & \text{if } k \text{ is odd},\\[4pt]
(k+1)(d-\ell) + \ell, & \text{if } k \text{ is even},
\end{cases}
\end{equation}
and the remaining \(k(d-\ell-1) + \ell - 1\) stabilizers each have weight \(2(d-\ell)\). For the additional repetition component of the outer layer, one stabilizer has weight \(d\), and the remaining \(d - 2\) stabilizers each have weight \(2(d-\ell)\). The average stabilizer weight is
\begin{equation}
    W_{\mathrm{both}}
    \approx
    \frac{
        (k_i+k(d-\ell))(2d-2\ell-1)+\cdots
    }{
        (k_i+k)(d-\ell-1)+\cdots
    }.
\end{equation}
In the limit \(k \gg d \gg 1\), this weight converges to \(W_{\mathrm{both}} \to 4\).

\subsubsection*{Protection}

As in the previous subsections, we verify that the concatenated construction attains distance \(d\) by analyzing all error patterns of weight at most \(d\). First, suppose that \(d' \le d\) \(X\)-type errors occur within the inner code of distance \(d\). Such events are uniquely identifiable by the inner layer, ensuring that all configurations with at most \(d\) \(X\) errors are detectable.

For \(Z\)-type errors, the analysis follows the same structure as before. If \(d'\) \(Z\) errors occur within any single repetition block, the effect is trivial when \(d'\) is even and corresponds to a single phase flip when \(d'\) is odd. In the worst case, therefore, up to \(d\) such phase flips may be introduced across the blocks. Take \( k_i = k(d-\ell) + \ell + (d - 1)\). If the inner code has parameters \([[k_i(d-\ell)+\ell,\, k_i,\, d]]\), it contains \(k_i + 1\) blocks, and the outer layer must protect against phase flips on all of them.

In this configuration, the outer overlapped-repetition code of distance \(d\) protects the \(k(d-\ell) + \ell\) blocks of length \(d-\ell\), while an additional outer repetition code of distance \(d\) protects the remaining \(d-1\) unshared blocks of length \(d-\ell\) together with the shared block of length \(\ell\). Since both outer codes have distance \(d\), all such phase-flip patterns remain detectable.

Together, these considerations establish that the concatenated construction achieves distance \(d\).



\subsection{Comparison}\label{subsec:compare}

Three integration strategies are summarized in Table~\ref{tab:compare}.

\begin{table}[ht]
\caption{Comparison of the resulting code parameters. For each construction, we set \(\ell = \lfloor d/2 \rfloor\), which optimizes the code rate. In all cases, the average stabilizer weight approaches \(4\) in the asymptotic regime \(k \gg d \gg 1\).}
\label{tab:compare}
\begin{tabularx}{\linewidth}{ c  c  c }
    \hline\hline
    Code & \hspace{0.4cm} & \([[n,k,d]]\) parameters \\
    \hline
    outer &  & \([[d(k\lceil d/2 \rceil + \lfloor d/2 \rfloor),\, k,\, d]]\) \\[1ex]
    inner &  & \([[((k+1)d-1)\lceil d/2 \rceil + \lfloor d/2 \rfloor,\, k,\, d]]\) \\[1ex]
    both  &  & \([[k\lceil d/2 \rceil^{2} + \lfloor d/2 \rfloor(\lceil d/2 \rceil + 1)+(d-1)\lceil d/2 \rceil,\, k,\, d]]\) \\[1ex]
    \hline\hline
\end{tabularx}
\end{table}
For a direct comparison, taking \(k=d=12\) in the construction that modifies both the inner and outer layers yields parameters \([[540,12,12]]\). This configuration requires more physical qubits than the \([[144,12,12]]\) BB code, but fewer than the \(1452\) (\(2028\)) qubits needed for \(12\) copies of the \([[121,1,11]]\) (\([[169,1,13]]\)) surface code. These values show that the overlapped-repetition Shor construction improves the code rate by a factor in the interval \([2.69,3.76]\) while maintaining a mean stabilizer weight of \(4\). Although this improvement is less dramatic than the \([10.08,14.08]\)-fold increase achieved by the BB code, at the cost of increasing the mean stabilizer weight to \(6\), the overlapped-repetition Shor family preserves a transparent syndrome-to-recovery mapping, thereby avoiding the performance-loss issues associated with iterative decoding.

\section{Further code constructions}\label{sec:further}

\subsection{LDPC design}

The structure identified in Sec.~\ref{subsec:both} suggests that the overlapped-repetition component of the outer layer may be replaced by any efficient classical \([n,k,d]\) code, at the cost of a potentially larger average stabilizer weight. If the chosen outer code is an LDPC code with maximum parity-check weight \(w\), then the corresponding quantum stabilizers have weight at most \(w(d-\ell)\) when \(w\) is even, and at most \(w(d-\ell)+\ell\) when \(w\) is odd.

In this setting, the inner layer must provide \(k_i = n + d - 1\) unshared blocks together with one shared block. The resulting quantum code therefore has parameters
\begin{equation}
[[\,k_i(d-\ell) + \ell,\, k,\, d\,]].
\end{equation}

Suppose the classical \([n,k,d]\) code is asymptotically good, meaning its rate approaches unity in the large-block-length limit. Under this assumption, the resulting quantum construction attains an asymptotic rate approaching \(1/(d-\ell)\), a substantial improvement over the typical \(1/d^2\) scaling.

\subsection{Collective coherent error adapted design}

In addition to stochastic Pauli noise, physical qubits are also subject to collective coherent (CC) errors~\cite{Bravyi2018, Debroy2018, HLRC22, Mrton2023, PhysRevA.90.062317, PhysRevA.93.042340, PhysRevA.111.052602, hwfz-c6vy, qzd2-b2mx}, arising from intrinsic single-qubit Hamiltonians. These errors can periodically reach high weight, and in most settings the precise time of occurrence is unknown, since the receiver may access the encoded state at an arbitrary moment.

A simple and effective method for immune CC errors is to employ constant-excitation encodings. This may be implemented by replacing each inner repetition block with a complement dual-rail map,
\[
(\overline{\ket{0}},\overline{\ket{1}})\ \mapsto\ (\ket{01},\ket{10}),
\]
so that every repetition block of even length maintains a fixed total excitation number. Concretely, the standard repetition basis
\[
(\ket{00\cdots 0},\ \ket{11\cdots 1})
\]
is replaced by
\[
\ket{01\cdots 01},\qquad \ket{10\cdots 10}.
\]

For the construction in Sec.~\ref{subsec:outer}, this substitution is valid whenever \(d\) is even. For the constructions in Sec.~\ref{subsec:inner} and Sec.~\ref{subsec:both}, both \(d-\ell\) and \(\ell\) must be even to ensure that every block has even length. This modification introduces no additional overhead and enhances robustness against CC noise, making it well suited for quantum-memory applications.

\subsection{Bosonic codes}

Interpreting all excitations within a (qubit) block of a repetition code as indistinguishable bosonic excitations of a single oscillator reproduces, in the simplest instances, the (extended) binomial codes~\cite{PhysRevX.6.031006,hwfz-c6vy}.
The same perspective applies directly to the overlapped-repetition constructions developed in this work, yielding a broad family of new bosonic codes obtained by bosonizing the corresponding qubit-based designs.

This approach expands the currently limited set of known bosonic examples, such as cat codes~\cite{PhysRevA.59.2631}, GKP codes~\cite{PhysRevA.64.012310}, and binomial codes, by providing a systematic route to generate additional constructions with tunable rate and distance.

\section{Conclusion}\label{sec:conclude}

We have shown that a simple modification of the standard Shor construction increases its code rate by a factor of four while preserving an asymptotic average stabilizer weight of \(4\). In addition, replacing the overlapped-repetition outer layer with a suitable LDPC code yields a family of quantum codes with asymptotic rate \(2/d\). We have also described a constant-excitation variant that provides protection against collective coherent errors at no additional overhead, as well as a bosonic generalization that extends the construction to oscillator-based codes.

The proposed family (with average stabilizer weight \(4\)) achieves a code rate intermediate between that of the rotated surface code (also with average stabilizer weight \(4\)) and the BB code (average stabilizer weight \(6\)), as detailed in Sec.~\ref{subsec:compare}. Moreover, decoding remains free of the performance degradation typically associated with iterative decoders. A systematic exploration of the trade-off between higher code rate and lower average stabilizer weight is left for future work.


\end{document}